# Interaction-Merged Motion Planning: Effectively Leveraging Diverse Motion Datasets for Robust Planning

### v1

@@ 1 Introduction
Motion planning [48, 77, 7, 76] is the final step in autonomous robot driving, focused on generating safe and efficient trajectory based on perception and prediction. With the availability of various trajectory datasets [6, 45, 34, 47, 4, 69, 38], data-driven approaches have become more popular due to their ability to adapt to diverse driving scenarios [56, 70, 20, 30, 21, 22].
@@ 2.2 Leveraging Diverse Motion Datasets
Motion datasets differ in data collection methods, environmental characteristics, dataset size, and diversity. As a result, imitation learning-based models trained for a specific domain often experience performance degradation when deployed in different domains [17, 16, 44]. To improve robustness in target domains, recent studies have investigated domain generalization at both the dataset and architecture levels. UniTraj [16] adopts a dataset-level approach by integrating multiple trajectory datasets into a unified dataset, demonstrating that incorporating diverse data enhances generalization. Ye et al. [74] enhance domain generalization by partitioning a single dataset based on vehicle motion characteristics and representing motion in a reference path-based frame. Wang et al. [61] introduce a module that refines domain-specific velocity and environmental characteristics, while Dong et al. [14] propose a Stepwise Attention Layer for domain alignment at each timestep. Model architecture-level approaches, such as those in [74, 61, 14, 58, 66, 46], are tailored to specific domains or models, limiting their applicability to general models. In contrast, IMMP operates independently of model architecture, making it adaptable across various planning models.
@@ 3.2 Model Merging and Task Vector
Model merging [72] is used to integrate the parameters of multiple network backbones, each trained on different tasks. This technique is primarily explored in language-image tasks [71, 54, 41] and image classification [23, 68]. Notably, Task Arithmetic [23] is a pivotal work in this domain, introducing the concept of the Task Vector, which enables simple parameter editing by adding or negating targeted information. The task vector is defined as follows:
@@ 4.3 Ablation Study
Table 3: Ablation study on checkpoint types and pool size used for merging. We evaluate the planning performance of IMMP when using the best checkpoints across all metrics, as well as intermediate epochs for planning (P) or forecaster (F) models. The fine-tuning performance of these checkpoints is also evaluated.
@@ 5 Conclusion
In this paper, we present the Interaction-Merged Motion Planning (IMMP) framework, which effectively leverages diverse motion planning datasets. By utilizing parameter checkpoints trained on different domains, IMMP addresses the challenges of domain imbalance and catastrophic forgetting that hinder conventional adaptation approaches. The proposed two-step process—interaction-conserving pre-merging and interaction-level merging—preserves agent behavior and interactions, resulting in a more adaptable planning model. Experimental results demonstrate that IMMP outperforms traditional approaches, achieving superior performance. Furthermore, IMMP seamlessly integrates with existing frameworks, enhancing its versatility and applicability across various motion planning tasks.

### v2

@@ 1 Introduction
Motion planning [54, 83, 7, 82] is the final step in autonomous robot driving, focused on generating safe and efficient trajectory based on perception and prediction. With the availability of various trajectory datasets [6, 51, 37, 53, 4, 75, 43, 24], data-driven approaches have become more popular due to their ability to adapt to diverse driving scenarios [62, 76, 20, 33, 21, 22, 25].
@@ 2.2 Leveraging Diverse Motion Datasets
Motion datasets differ in data collection methods, environmental characteristics, dataset size, and diversity. As a result, imitation learning-based models trained for a specific domain often experience performance degradation when deployed in different domains [17, 16, 49, 50]. To improve robustness in target domains, recent studies have investigated domain generalization at both the dataset and architecture levels. UniTraj [16] adopts a dataset-level approach by integrating multiple trajectory datasets into a unified dataset, demonstrating that incorporating diverse data enhances generalization. Ye et al. [80] enhance domain generalization by partitioning a single dataset based on vehicle motion characteristics and representing motion in a reference path-based frame. Wang et al. [67] introduce a module that refines domain-specific velocity and environmental characteristics, while Dong et al. [14] propose a Stepwise Attention Layer for domain alignment at each timestep. Model architecture-level approaches, such as those in [80, 67, 14, 64, 72, 52], are tailored to specific domains or models, limiting their applicability to general models. In contrast, IMMP operates independently of model architecture, making it adaptable across various planning models.
@@ 3.2 Model Merging and Task Vector
Model merging [78] is used to integrate the parameters of multiple network backbones, each trained on different tasks, and serves as an alternative to traditional multi-task learning [81, 39, 40, 26, 27] by avoiding joint training. This technique is primarily explored in language-image tasks [77, 60, 46] and image classification [23, 74]. Notably, Task Arithmetic [23] is a pivotal work in this domain, introducing the concept of the Task Vector, which enables simple parameter editing by adding or negating targeted information. The task vector is defined as follows:
@@ 4.3 Ablation Study
Table 3: Ablation study on checkpoint types and pool size for merging. IMMP planning performance with best checkpoints (across all metrics) and intermediate-epoch for planning (P) / forecaster (F) models; fine-tuning results included.
@@ 5 Conclusion
In this paper, we present the Interaction-Merged Motion Planning (IMMP), which effectively leverages diverse motion planning datasets. By utilizing parameter checkpoints trained on different domains, IMMP addresses domain imbalance and catastrophic forgetting that hinder conventional adaptation approaches. The proposed two-step process—interaction-conserving pre-merging and interaction-level merging—preserves agent behavior and interactions, resulting in a more adaptable planning model. Experimental results demonstrate that IMMP outperforms traditional approaches, achieving superior performance.
Acknowledgment
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (NRF2022R1A2B5B03002636), and by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00457882, AI Research Hub Project).